OpenStudy (kenialove):

What is the probability of tossing a penny and landing on heads three times in a row ?

OpenStudy (anonymous):

3/6

OpenStudy (steve816):

1/8

OpenStudy (anonymous):

you would toss this 3 times and its heads or tails so 3 out of 6

OpenStudy (steve816):

no that is not right @mathhelp1199

OpenStudy (anonymous):

i just did this about a month ago the number or favorable outcomes to number of total

OpenStudy (steve816):

it is \[\frac{ 1 }{ 2 }\times\frac{ 1 }{ 2 }\times \frac{ 1 }{ 2 }=\frac{ 1 }{ 8 }\]

OpenStudy (steve816):

@mathhelp1199 you clearly have not learned probability yet. The chance of a penny landing on heads 3 times in a row is 1/8

OpenStudy (anonymous):

well this makes no sense then cause i had the question on a test and got it right :l but oh well idc

OpenStudy (kenialove):

Can one of u show me how to find the answer at least ?

OpenStudy (steve816):

I already showed you look at the post above you.

OpenStudy (kenialove):

Oh okay thanks